EDITORIAL

Terry Goodman

As another school year winds down, the editors of MJMS want to thank you, our readers, for your support of the journal. We appreciate the many positive contributions you make to MJMS through the manuscripts, problems, problem solutions, and announcements that you send us. We also want to encourage you to communicate to us any suggestions you have for improving the journal. Finally, please let your colleagues know about MJMS and encourage them to subscribe and/or submit manuscripts, problems, and problem solutions.

I believe one of the reasons that MJMS has been successful is that we have a very dedicated and gifted editorial staff. Even though serving on our editorial staff is a “labor of love,” our editors have consistently done a wonderful job. One of our associate editors, Bill Grimes, is leaving his position since he is retiring from Central Missouri State University (the fish population of west central Missouri will definitely decrease with his retirement!). For those of us at MJMS who are from CMSU Bill has been a valued and trusted friend and colleague for many years. His work for our journal has consistently been of the highest quality and we will miss his thoughtful contribution.

We have recruited, however, a most able replacement for Bill. Vince Edmondson, a professor of Mathematics and Computer Science at CMSU will become our newest associate editor. Vince, like Bill, pays great attention to detail and quality. We are most pleased to have him join our editorial staff and anticipate that he will make a positive contribution to the quality of MJMS.